OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (campbell_st):

you know a = 10, n = 8 and term 8 so use the formula below \[T _{n} = ar^{n -1}\] \[781250 = 10 \times r^7\] solve for r when you have found r the sum of the x terms is \[S _{n} = \frac{a( r^n -1)}{r -1}\]

OpenStudy (anonymous):

Thank you :)